# Discrete Chi-Square Method beats Discrete Fourier Transform and other similar parametric time series analysis methods

### v3

@@ 4 Discussion
4.1 Existence ( $\mathrm{C}_{1}$)
[12] transformed the original function into the frequency domain. The modern DFT time series analysis method transforms the original data into the frequency domain and gives the best frequency for a pure sine model. [14] presented the least squares method, which minimises the differences between the data and the linear model. Our DCM does the same by testing a large number of linear models. The data spacing, even or uneven, is irrelevant for these models. For every chosen DCM model, the number of tested linear models is
$\binom{n_{\mathrm{L}}}{K_{1}}+(1+n_{\mathrm{B}})\times\binom{n_{\mathrm{S}}}{K_{1}}$, (46)
where the number of tested long and short search frequencies is $n_{\mathrm{L}}$ and $n_{\mathrm{S}}$, respectively. The number of bootstrap samples is $n_{\mathrm{B}}$. In other words, we solve this ill-posed problem by using brute computational force. If the data contain only zero mean white noise, the Gauß-Markov theorem ensures that a linear least squares fit solution exists for every tested frequency combination.
4.2 Uniqueness ( $\mathrm{C}_{2}$)
We make the most of the Gauß-Markov theorem. When the numerical values of the tested frequencies $\bm{\beta}_{I}$ (Equation 9) are fixed, the model becomes linear and the solution for the other free parameters $\bm{\beta}_{II}$ is unique (Equation 2.1). All possible $\bm{\beta}_{I}$ frequency combinations are tested (Equation 11). For every tested frequency combination $\bm{\beta}_{I}$, the linear model gives a unique value for the test statistic $z$ (Equations  15 and 16). From all tested frequency combinations, we select the best frequency combination $\bm{\beta}_{\mathrm{I,best}}$ which minimises $z$. The linear model for this best frequency combination $\bm{\beta}_{\mathrm{I,best}}$ gives unique values for the remaining other free parameters $\bm{\beta}_{\mathrm{II,best}}$. The only goal for the massive DCM search (Equation 46) is to find these unique initial free parameter values $\bm{\beta}_{\mathrm{initial}}=[\bm{\beta}_{\mathrm{I,best}},\bm{\beta}_{\mathrm{II,best}}]$ for the non-linear iteration that gives the unique final free parameter values $\bm{\beta}_{\mathrm{final}}$ (Equation 21).
We use the Fisher-test to compare many different non-linear DCM models against each other. The selection criterion for the best model is unique (Equation 40). The best DCM model is not necessarily the correct model, if this correct model is not among the compared models. The correct model must be able the predict the future and past data. We formulate the Predictability-test for alternative DCM models (Equation 3.10). The order of Fisher- and Predictability-tests can be reversed. However, the former uses all data, while the latter uses a subset of all data, the predictive data. In ideal cases, both tests identify the same best and correct model, like in Section 3.10.
4.3 Stability ( $\mathrm{C}_{3}$)
There are additional signatures of stability. The $z$ periodogram solution is unique for every tested $\bm{\beta}_{I}$ frequency combination. If these periodograms are continuous and their changes are not irregular (e.g., like in Figs. 1a-b), the DCM model solution is stable because it does not change by increasing the number of tested frequencies $n_{\mathrm{L}}$ and $n_{\mathrm{S}}$. Furthermore, the solutions for all seven simulated data sets converge when the $n$ and $\mathrm{SN}$} values increase (Tables 1 - 7). The different $n$ and $\mathrm{SN}$} combinations give the same stable DCM model solution.
We conclude that our computational statistical DCM model solution fulfils the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of the solution for a well-posed problem (Sections 4.1 - 4.3).
4.4 Performance
DCM does not suffer from the DFT limitations (Equations 27-30). The simulated complex data sets of Models 1-7 expose these DFT weaknesses, but in all these cases, DCM detects all periodic signals superimposed on an unknown trend (Sections 3.1 - 3.7). The correct signal(-s) and trend are detected. If there is enough $(n)$ very accurate $(\sigma_{i})$ data $(y_{i})$, DCM can predict the future or postdict the past by finding the correct DCM model for these data. We do admit that the $n$ and $\mathrm{SN}$} values of Models 2, 5 and 7 are extreme and unrealistic for most cases of real data. However, those model simulations are necessary for demonstrating that DCM can detect signals from short $\Delta T$ samples of high quality data. Those DCM signal detections depend only on the number of observations and their accuracy.
Our DCM computes the $R$ (Equation 13) or $\chi^{2}$ (Equation 14) values for a massive number of linear models (Equation 46). The Gauß-Markov theorem ensures that the model having the lowest $R$ or $\chi^{2}$ is eventually always found, and the result for the non-linear iteration is unique (Equation 21). We could have simulated more complex data sets by increasing the values of $K_{1}$ (number of signals), $K_{2}$ (signal shape) and $K_{3}$ (polynomial order of trend). DCM would eventually detect those signals and the trend when the sampled data sets get larger $(n)$ and/or the signal to noise ratio becomes higher $({\mathrm{SN}})$. In this sense, DCM has no limit for the complexity of the data sets. In the real world, it may be impossible to eliminate noise $({\mathrm{SN}})$, but it is possible increase the sample size $(n)$. Since the time span $(\Delta T)$ is irrelevant, the best solution is fast sampling of accurate data sets. For example, many observers could measure the same phenomenon simultaneously. We recommend such fast sampling combined to DCM parallel computation Python code analysis.
We want to draw attention to three aspects of DCM computational performance.
@@ 4.4 Performance
    Practical time series analysis applications, like digital signal processing, require fast computational algorithms [22]. Our computational statistical solution for the non-linear DCM model requires a lot of CPU. Yet, the CPU used is beside the point if a well-posed computational solution can be found to any challenging scientific non-linear problem. DCM proceeds through two stages. For every chosen alternative model, DCM performs a massive number of linear least squares fits (Equation 18) to find the unique initial free parameter values for the non-linear iteration (Equation 21). This first stage devours CPU. The second fast stage, the Fisher-test, then reveals the best model alternative (Equation 40).
@@
    Our frequency space symmetry idea dramatically reduces CPU (Equation 11). For example, were this symmetry not exploited, the four signal $K_{1}=4$ search would give $4!=24$ exactly the same solutions from four-dimensional tested grid in frequency space. This would cause unnecessary use of CPU in the preliminary testing all possible frequency combinations inside a tesseract (a four-dimensional cube). After one necessary iteration of Equation 21, and 23 unnecessary ones, the identification of the best frequency combination would be a daunting exercise.
@@
    The DCM signal detections depend only on the sample size $n$ and the signal to noise ratio $\mathrm{SN}$}, but not on the time span $\Delta T$. This dimensional effect makes sense. If the $t_{i}$ and $y_{i}$ data are inside a rectangle $\Delta T\times\Delta y=(t_{n}-t_{1})\times[{\mathrm{max}}(y_{i})-{\mathrm{min}}(y_{i})]$, then the $y_{i}/\sigma_{i}$ ratio reveals or hides the $g_{i}$ model details. Since the time span of data $\Delta T$ is irrelevant, DCM is an ideal forecasting method. We have been deluded into unnecessary waiting for the repetition of the signal(-s). DCM can perceive the future and the past much “earlier” (from shorter $\Delta T$) than previously thought, like in Equations 27 and 29. This should motivate fast sampling of large high quality data sets and their analysis using the DCM parallel Python code. The invention of computers has made it possible to use this computational statistical method, which is probably the simplest possible one.
The analytical solution for the non-linear DCM model is ill-posed (Equation 1). Regardless of the data set complexity, DCM can find a well-posed computational statistical solution for this problem (Equation 21). So far, we have compared DCM only to DFT. Both methods are parametric and operate in the frequency-domain. There are innumerable similar parametric time series analysis methods [e.g. 22, 15]. Each one of these other time series analysis methods suffers from their own particular application limitations. We list some of those application limitations below.
  1. 1.
    The required data spacing is even or uneven.
  2. 2.
    The number of signals is unknown.
  3. 3.
    The presence and shape of trend are unknown.
@@
    The degree of noise is unknown.
@@
    The sample window causes leakage.
@@
    The leakage weakens frequency resolution.
@@
    The signal frequencies are too close.
@@
    The signal periods exceed the data time span.
@@
    The signal shapes are not pure sinusoids.
@@
    The forecasts of non-linear models fail.
DCM does not suffer from any of these application limitations. This simple computational statistical method can fail if, and only if, the Gauß-Markov theorem is not valid. That is impossible.
@@
Author contribution Lauri Jetsu created the simulated data sets, performed the DCM and DFT analysis and wrote the manuscript.
@@
Consent for publication The author of this manuscript grants permission for the publication in Computational Statistics.

### v4

@@ 4 Discussion
DCM proceeds through two stages. It computes the $R$ (Equation 14) or $\chi^{2}$ (Equation 15) values for a massive number of LS fits (Equation 50). The Gauß-Markov theorem ensures that the DCM model having the lowest $R$ or $\chi^{2}$ is inevitably always found, and the result for the non-linear iteration is unique (Equation 22). This first stage devours CPU. The second fast stage, the Fisher-test, then reveals the best DCM model alternative (Equation 48). Practical time series analysis applications, like digital signal processing, require fast computational algorithms [25]. Long computation time is the main AL constraining DCM. For example, Jetsu (2025) computed about 4.5 million LS fits to search for four signals in the monthly sunspot data $(n=3287,K_{1}=4,K_{2}=1,K_{3}=0,n_{\mathrm{L}}=100,n_{\mathrm{S}}=30,n_{\mathrm{B}}=20)$. This required several months of CPU. The parallel Python code computations took a few days. Nevertheless, the CPU used is beside the point if a well-posed computational solution can be found to any challenging scientific ill-posed problem.
The DCM computation time AL constraint is amply compensated in real observations because there is no need to wait for the repetition of the signal(-s). The WD-effect is the spearhead of DCM because the sample window $(\Delta T)$ can be infinitesimally short. Fast accurate observations are the best approach. If the noise $(\sigma\equiv{\mathrm{SN}})$ can not be eliminated it is always possible to increase the sample size $(n)$.
It is time to summarise why the ALs of other frequency-domain parametric time series analysis methods (Section 1) do not constrain DCM.
@@
    Data errors (level of noise) are unknown.
    • DCM can solve both alternatives: errors $\sigma_{i}$ known or unknown (Equations 16, 17, 46 and 47). All DCM model parameter solutions in Tables 1-7 converge to the correct values when the simulated data $n$ and/or $\mathrm{SN}$} increase ($\sigma_{i}$ decrease). Due to this WD-effect, sufficient noise reduction always leads to the correct model detection.
@@
    Data error information is not utilised.
    • DCM utilises this information. If the errors $\sigma_{i}$ are known, DCM performs weighted LS fits which minimise the model $\chi^{2}$ for every tested frequency combination (Equation 15). This gives the DCM test statistic $z$ (Equation 17). The Fisher-test utilises the $\chi^{2}$ values of all alternative tested DCM models to identify the best DCM model (Equation 47).
@@
    Data must be evenly spaced.
    • DCM performance is independent of data spacing. Even or uneven spacing is irrelevant for all $n_{\mathrm{Lin}}$ LS fits (Equation 50). However, we do admit that long gaps can mislead even these LS fits.
@@
    Model parameter errors are unknown.
    • DCM gives error estimates for the model parameters of Equations 23-29. For non-linear models, such as the DCM model, the analytical solution for the model parameter errors is a highly complex effort [12]. DCM solves these error estimates using the computational statistical bootstrap technique (Equation 31).
@@
    Model and forecast errors are unknown.
    • The computational bootstrap technique (Equation 31) gives $n_{B}$ estimates for $\bm{\beta}_{J}$, $t_{\mathrm{mid,J}}$ and $\Delta T_{J}$, where $J=1,2,...n_{B}$. These estimates give $n_{B}$ values for $g(t)$ at any time $t$ inside and outside the sample window $\Delta T$. The standard deviation of these $n_{B}$ values gives the error limits $\sigma_{g(t)}$ in Equation 30. The $n_{B}$ estimates for $\bm{\beta}_{J}$, $t_{\mathrm{mid,J}}$ and $\Delta T_{J}$ also give the error limits for the functions $h(t),h_{i}(t)$ and $p(t)$ in the DCM model $g(t)$.
@@
    Sample window is shorter than signal period(s).
    • DCM solves this problem. The sample window is shorter than the period(s) in Model 1, 2, 5 and 7 time series simulations. Regardless of this, all model parameter estimates converge to correct values when $n$ and/or $\mathrm{SN}$} increase (Tables 1, 2, 5 and 7). Due to this WD-effect, the sample window $\Delta T$ value is irrelevant for DCM. It can perceive the future and the past much “earlier” (from shorter $\Delta T$) than has been previously thought, like in Equations 35 and 37.
@@
    Presence and shape of trend are unknown.
    • DCM can test any $K_{1}$, $K_{2}$ and $K_{3}$ model combination. The Fisher-test reveals the combination of the best DCM model. The Forecast-test can double-check this result. These tests give the correct $K_{3}$ trend order. Due to the WD-effect, the polynomial trend coefficients $M_{k}$ converge to correct values in all simulated time series (Sections 3.1-3.7). The trend is absent (stationary time series) if $K_{3}=-1\equiv$ $p(t)=0$ or $K_{3}=0\equiv$ $p(t)=M_{0}=\text{constant}$ (Equations 4 and 5).
@@
    Sample window causes leakage.
    • Leakage does not constrain DCM because the sample window $\Delta T$ has no effect on the LS fits (Equation 50: $n_{\mathrm{Lin}}$). All frequencies converge to exactly correct values when the $n$ and/or $\mathrm{SN}$} values of simulated time series increase (Tables 1-7). This confirms the absence of leakage. Due to the WD-effect, the performance of DCM does not depend on the sample window $\Delta T$.
@@
    Leakage weakens frequency resolution.
    • There is no leakage because the sample window is irrelevant (AL8). The DCM frequency resolution is limited only by the sample size $n$ and the data accuracy $\sigma$. The WD-effect ensures that all frequency detections for the simulated time series converge to exactly correct values (Tables 1-7).
@@
    Signal shapes are not pure sines.
    • The DCM always finds the correct $K_{1}$, $K_{2}$ and $K_{3}$ values (AL7). The $K_{2}$ value determines the signal shape. The $K_{2}=1$ pure sine shape is the simplest. Higher $K_{2}$ values allow the DCM modelling of more complex shapes. We demonstrate the $K_{2}=2$ double wave signal detections in Sections 3.6 and 3.7.
  11. 11.
    Number of signals is unknown.
[…]
  12. 12.
    Correct model alternative is unknown.
    • The values of $K_{1}$, $K_{2}$ and $K_{3}$ determine the correct model (AL7). The $(n-\eta_{2}-1)/(\eta_{2}-\eta_{1})$ penalty terms prevent overfitting when the Fisher-test is used to identify the best DCM model from all tested DCM models (Equations 46 and 47). This best model is also the correct model if it passes the Forecast-test (Equation 3.12). The WD-effect ensures that better data can inevitably reveal the correct model.
  13. 13.
    Signal significances are unknown.
    • The Fisher-test gives the $Q_{\mathrm{F}}$ critical levels (signal significances) for all detected signals (Sections 3.9 and 3.10). This $Q_{\mathrm{F}}$ is the probability of falsely rejecting the $H_{0}$ hypothesis when it is in fact true. Thus, the $Q_{\mathrm{F}}$ value represents the probability of false signal detection. This detection is absolutely certain in the $Q_{\mathrm{F}}<\!10^{-16}$ cases. Our pre-assigned significance level for signal detection is $\gamma=0.001$ (Equation 48).
  14. 14.
    Model solution is ill-posed.
    • The analytical solution for the non-linear DCM model is ill-posed (Equation 1). We show that there exists a well-posed computational solution (Section 3.11). The WD-effect ensures that this solution can always be found by increasing the sample size $(n)$ and/or the signal to noise ratio $({\mathrm{SN}})$, regardless of the time series complexity ($K_{1}$, $K_{2}$ and $K_{3}$ combination). The main DCM constraint is that the solutions for more complex models require more computation time when the number of signals increases (Equation 50).
  15. 15.
    Complex non-linear model forecasts fail.
    • Our computational solution for the non-linear DCM model is well-posed (Section 3.11). The forecast of this well-posed solution can be double-checked (Section 3.12). First, the Fisher-test identifies the best DCM model from all tested DCM models (Section 3.9). Then, the Forecast-test reveals if this best DCM model is also the correct one (Equation 3.12). Again, the WD-effect ensures that the correct model and forecast can be found for any complex time series ($K_{1}$, $K_{2}$ and $K_{3}$ combination) when $n$ and/or $\mathrm{SN}$} increase. We conclude that the DCM rises to meet the challenge of “forecasting the evolution of complex systems” [7].
The DCM turns things upside down. There would be no need for time series analysis, if the correct frequencies were already known. The tested frequencies are already known. DCM does not search for unknown frequencies, it just tests known frequencies. DCM tests all possible frequency combinations. The LS fit for every frequency combination gives unique $R$ and $\chi^{2}$ values. The best frequency combination minimises $R$ or $\chi^{2}$. This gives the unique initial free parameter $\bm{\beta}$ values for the non-linear iteration. The frequencies are never unknown in this process. The Gauss-Markov theorem, the well-posed computational model solution and the revolutionary WD-effect make absolutely sure that DCM succeeds for any number of signals $(K_{1})$, all signal shapes $(K_{2})$ and every polynomial trend $(K_{3})$. The Fisher-test and the Forecast-test can double-check that the DCM model solution is correct. The DCM analysis of intensive, large and accurate time series can see through time: a glimpse of the future and the past.
Our DCM is a remarkable method because it outperforms the DFT. Gauß (1777–1855) developed an early form of the Fast Fourier Transform (FFT) algorithm for astronomical purposes, but did he not publish it. Perhaps our DCM should be re-named the Slow Gauß Transform (SGT).
@@
Author contribution Lauri Jetsu performed this research and wrote the manuscript.
@@
Consent for publication The author of this manuscript grants permission for the publication in Journal of Statistical Theory and Practice.